POTENTIAL REWARDS FROM THE BEGINNING TO THE END OF A SPECTROSCOPIC CAREER DERIVED FROM ATTENDANCE AT A GRADUATE-STUDENT AND POSTDOCTORAL-FELLOW ORI-ENTED SYMPOSIUM

\underline{JON T. HOUGEN}, OPTICAL TECHNOLOGY DIVISION, NATIONAL INSTITUTE OF STAN-DARDS AND TECHNOLOGY, GAITHERSBURG, MD 20899-8441, USA.

In this talk we will consider some generalizations (based primarily on the authors observations and experiences at 52 of the past 65 Columbus meetings) concerning the benefits and pleasures of OSU Symposium attendance to a molecular spectroscopist at each stage of his or her career, i.e., undergrad, grad student, post doc, new staff member, mature staff member, and retiree. In particular, we will examine the three topics of: (i) what type of people frequently attend this conference, (ii) what do they gain in terms of things received from the conference, and (iii) what do they gain by contributing to the conference.